User name Punjabi
Log entry time 23:58:43 on March28,2000
Entry number 37145
16:15 MCC is trying to get beam back to do Arc energy measurement. TCarbon target is in. Raster is off.
17:40, arc measurement done (Doug Higinbothom)
18:00 begin harp scan, 160 umx230 um
19:25 accept beam with harp scan results 59 um x 90 um .
enabled raster, moved target to 3He.
beam on target, 5 uA, spot size 4x4 mm^2
start run in 32/23 kinematics at 19:40
finished run # 2993 at 19:48, 805 K events.
19:48 Started configuration change for kinematics # 23.
The Hadron arm is moved to 25.52 deg. The Hadron arm
momentum is set at 1.295 GeV/c.
start 2994 at 20:20; event rate is 500 Hz. Got 1 M events. Good run.
start 2995 at 20:59
21:12 energy lock now on Hall A
21:20 beam off due to RF problems. Stop Run 2995. 150 K events.
21:50 MCC called. Water leak in injector. Will take several hours to cleanup.
No beam expected before midnight.